Please show full work detail.

How much work must be done on 1.2 moles of an ideal gas at 430 K to reduce its volume by a factor of 0.19 (i.e., Vf = 0.19 Vi) at constant temperature?

Give your answer in Joules, as a positive number (the external work must be positive to compress a gas), to three signficant digits.

Your Answer:
Answer #1

workdone = nRTlnV2/V1 = 1.2*8.314*430*ln0.19 = -7124.58 J or 7.125 kJ